# The Trojan Knowledge: Bypassing Commercial LLM Guardrails via Harmless Prompt Weaving and Adaptive Tree Search

### v2

@@ 1 Introduction
Current jailbreak research, however, fundamentally overlooks a more intrinsic threat: the knowledge within target LLMs is not atomic but inherently interconnected [wei2025llms, wu2024evaluating]. Consequently, restricted facts can be reconstructed through a sequence of related sub-facts. While alignment systems may successfully block direct inquiries for harmful information, malicious objectives can still be realized by sequentially “weaving” together queries for decomposed sub-facts. Operationalizing this fundamental vulnerability requires adhering to three core principles. First, the attack must be assembled from a sequence of locally innocuous queries that deliberately exploit knowledge correlations; these interactions appear benign in isolation yet become informative when combined. Second, decomposition must rely on the target model’s internal knowledge; as attackers typically seek information they lack, the strategy should be to leverage the target model’s responses to bridge the expertise gap rather than relying on the attacker’s limited priors. Third, the process demands adaptive and dynamic exploration. By utilizing the target’s responses as guidance, the attack can navigate multiple reasoning paths (e.g., pivoting between distinct TNT synthesis routes like the traditional German process vs. the US mixed-acid process), ensuring that exploration continues even if a specific path is obstructed.
@@ 2 Related Work
Prompt-optimization methods formulate jailbreaking as a strategic search within the prompt space, aiming to directly elicit harmful outputs from target LLMs. Early approaches such as GCG [zou2023universal] and AutoDAN [liu2023autodan] employ gradient-based or genetic algorithms to generate adversarial inputs, but often produce unnatural artifacts easily intercepted by modern safety filters [zeng2024johnny]. Subsequent works have expanded this search approach through diverse mechanisms: quality-diversity and mutation pipelines (ReNeLLM [ding2023wolf], FERRET [Pala-ferret]), competitive fuzzing and random-search strategies [yu2023gptfuzzer, yao2023fuzzllm, zhou2025tempest0], and obfuscation or cross-lingual techniques exploiting robustness gaps [husain2025alphabet, yong2024lowresourcelanguagesjailbreakgpt4]. Persuasion-aware methods like PAP [zeng2024johnny] further augment attacks by integrating rhetorical variations. More recently, the field has shifted toward iterative and agentic refinement: systems such as GOAT [pavlova2024automatedredteaminggoat], Strategize-Adapt [chen2025strategizegloballyadaptlocally], and Chain-of-Attack [yang2024chainattacksemanticdrivencontextual] leverage target or judge feedback for adaptation, while others employ DRL-guided optimization [chen2025llmmeetsdrladvancing] or judge-guided mechanisms to identify interpretable jailbreaks [chao2025jailbreaking, inan2023llama]. Adaptive methods such as [qi2025majic, yu2025adaptive, andriushchenko2024jailbreaking] iteratively adjust and restructure prompts, but still operate primarily as prompt-optimization approaches within the prompt space. Additional methods explored prompt optimization strategies such as learning reusable strategies, parallel searching, etc. [li2025eliciting, chowdhury2025jailbreaking, mehrotra2024treeattacksjailbreakingblackbox, wu-etal-2025-monte, zhou2025tempestautonomousmultiturnjailbreaking]. Despite these advances, such methods remain fundamentally prompt-centric: they repeatedly query the target to optimize single-shot prompts or templates that explicitly encode malicious intent, rendering them increasingly vulnerable to pattern-matching defenses [muhaimin2025helpinglargelanguagemodels] and stronger alignment training [Zhang2023DefendingLLA, Zong2024SafetyFAA], while incurring substantial query costs.
@@
Decomposition-based jailbreaks seek to evade intent detectors by splitting a harmful objective into a sequence of seemingly benign sub-queries whose outputs can implicitly reconstruct the target goal. However, existing approaches generally rely on a static, up-front decomposition plan produced by a dedicated decomposer [wahreus2025prompt, srivastav2025safe]. Such plans do not adapt to target-model feedback during execution and often require substantial prior knowledge about the domain or model behavior. Other lines of work [xu2024redagentredteaminglarge, zhou2025autoredteamerautonomousredteaming, chen2024pandora] employ multi-agent frameworks to incorporate memory, reflection, or strategy selection, or leverage implicit-reference attacks [wu2024knowimsayingjailbreak] to obscure intent. Nevertheless, these systems still lack dynamic, feedback-driven decomposition, limiting their ability to probe nuanced or domain-specific safety vulnerabilities.
@@ 3.1 Problem Formulation
Under this framework, any jailbreak mechanism $\mathcal{A}$ represents a strategy for discovering a path $\mathcal{P}=(v_{0}\rightarrow\cdots\rightarrow v_{\mathrm{ans}})$. Critically, within a knowledge-rich target LLM, $\mathcal{G}$ is densely connected, offering multiple distinct paths to reach $v_{\mathrm{ans}}$ (e.g., different chemical synthesis processes for TNT). Based on this, our CKA-Agent operationalizes this latent DAG by “unrolling” potential paths through iterative dynamic interaction, transforming abstract graph exploration into the concrete, tractable tree search algorithm described below.
@@ 3.3 The Adaptive Branching Search Algorithm
(c) Execution and Hybrid Evaluation. To replace costly rollouts in traditional MCTS, the algorithm executes each generated sub-query $x^{(j)}$ against the target model to obtain $y^{(j)}$. The resulting pair is immediately scored by the Hybrid Evaluator, which linearly two metrics: an Introspection Score (assessing logical coherence and goal relevance) and a Target Feedback Score (capturing the information gain from the target LLM response and penalizing refusals). A linear combination of these scores becomes the initial $f_{v}$ for the newly created child node.
@@ 4.1 Settings
We compare CKA-Agent against a diverse set of representative jailbreak methods spanning both POA and DA approaches. The first group focuses on optimizing or refining prompts to elicit restricted responses. Vanilla directly queries the target model with the original harmful prompt, measuring inherent model robustness. AutoDAN [liu2023autodan] employs a hierarchical genetic algorithm with sentence-level momentum word scoring and paragraph-level elitist selection to generate jailbreak prompts. PAIR [chao2025jailbreaking] (Prompt Automatic Iterative Refinement) uses an attacker LLM to iteratively refine prompts based on judge feedback, discovering interpretable jailbreaks in approximately 20 queries without requiring gradient access. PAP [zeng2024johnny] (Persuasive Adversarial Prompts) leverages social science persuasion taxonomies to rewrite harmful queries; we evaluate the five most effective techniques: Logical Appeal, Authority Endorsement, Misrepresentation, Evidence-based Persuasion, and Expert Endorsement. ActorBreaker [ren2024derailyourselfmultiturnllm] and X-Teaming [rahman2025xteamingmultiturnjailbreaksdefenses] further extend this paradigm to multi-turn settings, where attacker agents iteratively refine malicious queries through conversational interactions. The second group, DA methods, instead split the harmful objective into sub-queries. Multi-Agent Jailbreak [srivastav2025safe] partitions a harmful query into multiple seemingly innocuous sub-queries, obtains the target model’s responses to each, and aggregates these responses to construct the final harmful output.
@@ 4.2 Main Results
Figure 4: Comparison of jailbreak methods across multiple defenses (Target LLM: Gemini-2.5-Flash). The $x$-axis corresponds to methods; each group contains four bars for No Defense, LLM Guard, Rephrasing, and Perturbation. The $y$-axis reports Attack Success Rate $(\mathrm{FS}\uparrow)$.
Existing Prompt- and Representation-Level Defenses Largely Fail Against Decomposition-Based Jailbreaks. We evaluate three families of external defenses that protect a target model without modifying its parameters at test time: (i) detection-based filters, for which we use Llama Guard-3 [dubey2024llama3herdmodels]; (ii) mutation-based defenses, which rephrase the input with a neutral LLM or apply light character/whitespace perturbations; and (iii) a representation-based defenses, such as Circuit Breaker [zou2024improving], which finetunes a target model at the level of internal representations to directly suppress harmful outputs, serving as an alternative to refusal or adversarial training. Since we cannot finetune the target model in our black-box setting, we instead apply the official finetuned Circuit Breaker model from [zou2024improving] in the same way as Llama Guard to defend against all attack methods. We omit perplexity-based filtering [alon2023detecting] because it requires logits unavailable in our black-box setting. As shown in Fig. 4, detection-based defenses, and especially the representation-based Circuit Breaker, substantially reduce the success of single-shot, prompt-over-optimization attacks such as AutoDAN, PAIR, PAP, and ActorBreaker, indicating that conspicuously adversarial prompts can often be intercepted. However, mutation-based defenses offer only modest additional protection: modern LLMs are highly tolerant to small lexical changes, with notable degradation mainly for pattern-sensitive methods like AutoDAN, and all prompt-level defenses struggle against multi-turn, decomposition-based attacks (X-Teaming, Multi-Agent Jailbreak, CKA-Agent). For these methods, individual sub-queries or single-turn messages may appear benign and are distributed across multiple turns or agents, so a defense that operates on isolated prompts is inherently limited. We also observe slight increases in attack success for certain methods, potentially because rephrasing makes some unnatural attack patterns less detectable, or because refusals from Llama Guard encourage attack algorithms to explore more attempts and strategies, which again underscores the limited effectiveness of current defenses. Among all defense methods, Circuit Breaker remains the most effective against our CKA-Agent. This may largely benefits from its training approach, which aims to remap internal model representations associated with harmful processes into incoherent or refusal representations. However, even though such representation-level control is more effective than pure prompt-level defenses, existing mechanisms remain far from satisfactory, likely due to their limited ability to perform multi-turn reasoning and aggregate signals to infer harmful intent. We further discuss potential defense directions in Sec. 4.5.
@@ 4.3 Alignment Between Human and LLM Judgments
To assess the calibration of our LLM-as-Judge ASR metric, we recruited ten doctoral researchers specializing in Electrical Engineering and Computer Science (EECS) to conduct a human evaluation study. The study employed a between-subjects design on 40 randomly sampled prompt–response pairs (10 per category: FS, PS, V, R), with the evaluation session lasting two hours per annotator. Five annotators evaluated the pairs using only the prompt and response (No Reasoning Provided), while a distinct group of five evaluators assessed the same items with access to the judge model’s reasoning (With Reasoning Provided). As shown in Fig. 7, alignment with the LLM judge increases substantially, from a Spearman correlation of 0.52 in Condition 1 to 0.90 in Condition 2. Crucially, the high variance observed in the No Reasoning setting aligns with our dataset design, which intentionally targets high-stakes domains requiring specialized knowledge (e.g., Chemistry, Biology). Since our evaluators possess deep expertise in EECS but lack specific domain knowledge in these external fields, they exhibited difficulty and a leniency bias when assessing the technical actionability of the responses without guidance. Providing the judge model’s analytical reasoning effectively bridges this domain gap, leading human annotators to evaluations that closely track the LLM’s assessments. These findings confirm that when supported by relevant domain reasoning, the judge model is well-calibrated and aligns closely with expert human judgment.
@@ 4.5 Evaluating Malicious Intent Detection Capabilities via Single-Session Decomposition
We further posit that CKA-Agent serves as a critical diagnostic framework for studying a fundamental limitation in current alignment: the inability to infer malicious intent when it is semantically distributed across a sequence of innocuous turns. While prior decomposition studies often leverage cross-session attacks to evade history tracking mechanisms, such settings deliberately obscure the model’s inherent multi-turn reasoning capabilities by introducing memory loss as a confounding variable. To rigorously evaluate whether aligned models can detect malicious users when given maximal visibility, we propose the single-session CKA-Agent-Branch as a stress test for intent aggregation. In this setting, each sub-query is submitted alongside the entire conversation history of the current branch. This effectively simulates a context-aware defense where the model has “perfect memory,” allowing us to isolate and assess the model’s cognitive ability to connect benign sub-tasks into a harmful whole. Table 3 validates the utility of this framework. Even with full history visibility, the attack maintains Full Success rates exceeding 92% across most target models. For instance, while GPT-oss exhibits marginally greater resistance compared to the cross-session setting, its consistently high failure rate exposes a structural deficiency: current alignment processes fine-tune models to detect atomic harmful prompts, but fail to instill the capability to aggregate intent over extended dialogs. This observation echoes the previous studies [zhou2024speak, liu2024lost, laban2025llms]. Consequently, we argue that CKA-Agent-Branch provides a robust testbed for future research, distinguishing between simple pattern matching and genuine, long-horizon intent understanding.
@@
Table 3: Performance comparison between CKA-Agent and CKA-Agent-Branch across different models and datasets. Results are reported for Gemini-2.5-Flash, Gemini-2.5-Pro, GPT-oss, and Claude-Haiku-4.5 on HarmBench and StrongREJECT datasets. FS: Full Success rate, PS: Partial Success rate, V: Violation rate, R: Rejection rate. Best results are highlighted in bold.

### v3

@@ 1 Introduction
Current jailbreak research, however, overlooks a more intrinsic threat: the knowledge within target LLMs is not atomic but inherently interconnected [wei2025llms, wu2024evaluating]. Consequently, restricted facts can be reconstructed through a sequence of related sub-facts. While alignment systems may successfully block direct inquiries for harmful information, malicious objectives can still be realized by sequentially “weaving” together queries for decomposed sub-facts. Operationalizing this fundamental vulnerability requires adhering to three core principles. Principle I: the attack must be assembled from a sequence of locally innocuous queries that deliberately exploit knowledge correlations; these interactions appear benign in isolation yet become informative when combined. Principle II: decomposition must rely on the target model’s internal knowledge; as attackers typically seek information they lack, the strategy should be to leverage the target model’s responses to bridge the expertise gap rather than relying on the attacker’s limited priors. Principle III: the process demands adaptive and dynamic exploration. By utilizing the target’s responses as guidance, the attack can navigate multiple reasoning paths (e.g., pivoting between TNT synthesis routes like the traditional German process vs. the US mixed-acid process), ensuring that exploration continues even if a specific path is obstructed.
@@ 2 Related Work
Prompt-optimization methods formulate jailbreaking as a strategic search within the prompt space, aiming to directly elicit harmful outputs from target LLMs. Early approaches such as GCG [zou2023universal] and AutoDAN [liu2023autodan] employ gradient-based or genetic algorithms to generate adversarial inputs, but often produce unnatural artifacts easily intercepted by modern safety filters [zeng2024johnny]. Subsequent works have expanded this search approach through diverse mechanisms: quality-diversity and mutation pipelines (ReNeLLM [ding2023wolf], FERRET [Pala-ferret]), competitive fuzzing and random-search strategies [yu2023gptfuzzer, yao2023fuzzllm, zhou2025tempest0], and obfuscation or cross-lingual techniques exploiting robustness gaps [husain2025alphabet, yong2024lowresourcelanguagesjailbreakgpt4]. Persuasion-aware methods like PAP [zeng2024johnny] further augment attacks by integrating rhetorical variations. More recently, the field has shifted toward iterative and agentic refinement: systems such as GOAT [pavlova2024automatedredteaminggoat], Strategize-Adapt [chen2025strategizegloballyadaptlocally], and Chain-of-Attack [yang2024chainattacksemanticdrivencontextual] leverage target or judge feedback for adaptation, while others employ DRL-guided optimization [chen2025llmmeetsdrladvancing], based on human-provided red-teaming principles to orchestrate prompts [xiong2025cop],or judge-guided mechanisms to identify interpretable jailbreaks [chao2025jailbreaking, inan2023llama]. Adaptive methods such as [qi2025majic, yu2025adaptive, andriushchenko2024jailbreaking] iteratively adjust and restructure prompts, but still operate primarily as prompt-optimization approaches within the prompt space. Additional methods explored prompt optimization strategies such as learning reusable strategies, parallel searching, etc. [li2025eliciting, chowdhury2025jailbreaking, wu-etal-2025-monte]. Noticeably, even though previous methods [mehrotra2024treeattacksjailbreakingblackbox, zhou2025tempestautonomousmultiturnjailbreaking] attempts to utilize tree structures in exploring, they inherently conduct prompt optimizations, where leafs maybe an updated harmful prompt [mehrotra2024treeattacksjailbreakingblackbox] or harmful conversations [zhou2025tempestautonomousmultiturnjailbreaking]. Despite these advances, such methods remain fundamentally prompt-centric: they repeatedly query the target to optimize single-shot prompts or templates that explicitly encode malicious intent, rendering them increasingly vulnerable to pattern-matching defenses [muhaimin2025helpinglargelanguagemodels] and stronger alignment training [Zhang2023DefendingLLA, Zong2024SafetyFAA], while incurring substantial query costs.
@@
Decomposition-based jailbreaks seek to evade intent detectors by splitting a harmful objective into a sequence of seemingly benign sub-queries whose outputs can implicitly reconstruct the target goal. However, existing approaches generally rely on a static, up-front decomposition plan produced by a dedicated decomposer [wahreus2025prompt, srivastav2025safe, brown2025benchmarkingmisusemitigationcovert]. Such plans do not adapt to target-model feedback during execution and often require substantial prior knowledge about the domain or model behavior. Similarly, prior work [zhang-etal-2025-damon] leverages tree search for decomposition but retains the attacker as a static planner, strategically revising queries upon refusal. Other lines of work [xu2024redagentredteaminglarge, zhou2025autoredteamerautonomousredteaming, chen2024pandora] employ multi-agent frameworks to incorporate memory, reflection, or strategy selection, or leverage implicit-reference attacks [wu2024knowimsayingjailbreak] to obscure intent. Nevertheless, these systems still lack dynamic, feedback-driven decomposition, limiting their ability to probe nuanced or domain-specific safety vulnerabilities.
@@ 3.1 Problem Formulation
Under this framework, any jailbreak mechanism $\mathcal{A}$ corresponds to discovering a path $\mathcal{P}=(v_{0}\rightarrow\cdots\rightarrow v_{\mathrm{ans}})$ within the latent reasoning DAG $\mathcal{G}$. Critically, within a knowledge-rich target LLM, $\mathcal{G}$ is densely connected, offering multiple distinct paths to reach $v_{\mathrm{ans}}$ (e.g., different chemical synthesis processes for TNT). However, because the adversary can only interact with $\mathcal{M}_{\mathrm{target}}$ through black-box queries, this latent DAG is not directly observable. Instead, each interaction conditionally expands a concrete reasoning trajectory based on previously elicited responses, implicitly unrolling the latent DAG into a tree-structured search space. As a result, the jailbreak process necessarily unfolds as a branching tree. This formulation naturally supports progressive decomposition, hypothesis branching, and selective exploration, properties that have long underpinned effective solutions in planning and search problems. Motivated by this insight, CKA-Agent explicitly operationalizes jailbreaking as a structured tree search process, systematically expanding and prioritizing reasoning branches through iterative interaction, thereby transforming abstract latent graph traversal into a tractable and controllable search procedure.
@@ 3.3 The Adaptive Branching Search Algorithm
(c) Execution and Hybrid Evaluation. To replace costly rollouts in traditional MCTS, the algorithm executes each generated sub-query $x^{(j)}$ against the target model to obtain $y^{(j)}$. The resulting pair is immediately scored by the Hybrid Evaluator, which linearly combines two metrics: an Introspection Score (assessing logical coherence and goal relevance) and a Target Feedback Score (capturing the information gain from the target LLM response and penalizing refusals). A linear combination of these scores becomes the initial $f_{v}$ for the newly created child node.
@@ 4.1 Settings
We compare CKA-Agent against a diverse set of representative jailbreak methods spanning both POA and DA approaches. The first group focuses on optimizing or refining prompts to elicit restricted responses. Vanilla directly queries the target model with the original harmful prompt, measuring inherent model robustness. AutoDAN [liu2023autodan] employs a hierarchical genetic algorithm with sentence-level momentum word scoring and paragraph-level elitist selection to generate jailbreak prompts. PAIR [chao2025jailbreaking] (Prompt Automatic Iterative Refinement) uses an attacker LLM to iteratively refine prompts based on judge feedback, discovering interpretable jailbreaks in approximately 20 queries without requiring gradient access. PAP [zeng2024johnny] (Persuasive Adversarial Prompts) leverages social science persuasion taxonomies to rewrite harmful queries; we evaluate the five most effective techniques: Logical Appeal, Authority Endorsement, Misrepresentation, Evidence-based Persuasion, and Expert Endorsement. TAP [mehrotra2024treeattacksjailbreakingblackbox] utilizes tree structures in exploring and optimizing jailbreaking prompts. ActorBreaker [ren2024derailyourselfmultiturnllm] and X-Teaming [rahman2025xteamingmultiturnjailbreaksdefenses] further extend this paradigm to multi-turn settings, where attacker agents iteratively refine malicious queries through conversational interactions. The second group, DA methods, instead split the harmful objective into sub-queries. Multi-Agent Jailbreak [srivastav2025safe] partitions a harmful query into multiple seemingly innocuous sub-queries, obtains the target model’s responses to each, and aggregates these responses to construct the final harmful output.
@@ 4.2 Main Results
Figure 4: Comparison of jailbreak methods across multiple defenses (Target LLM: Gemini-2.5-Flash; Dataset: HarmBench). The $x$-axis corresponds to methods; each group contains four bars for No Defense, LLM Guard, Rephrasing, and Perturbation. The $y$-axis reports Attack Success Rate $(\mathrm{FS}\uparrow)$.
Existing Prompt- and Representation-Level Defenses Largely Fail Against Decomposition-Based Jailbreaks. We evaluate three categories of inference-time defenses: (i) detection-based filters (Llama Guard-3 [dubey2024llama3herdmodels]); (ii) mutation-based defenses, which employ rephrasing or character-level perturbations; and (iii) representation-based defenses, such as Circuit Breaker [zou2024improving], which suppresses harmful outputs by intervening at the internal representation level. Given our black-box setting, we utilize the official Circuit Breaker model as a standalone guardrail and omit perplexity-based filtering [alon2023detecting] due to the lack of logit access. As shown in Fig. 4, detection-based defenses and Circuit Breaker significantly reduce the success of optimization-based attacks such as AutoDAN, PAIR, PAP, ActorBreaker and TAP, indicating that conspicuously adversarial prompts are often interceptable. Conversely, mutation-based defenses offer only modest protection because modern LLMs are highly robust to minor lexical changes. Crucially, all prompt-level defenses struggle against multi-turn, decomposition-based attacks including X-Teaming, Multi-Agent Jailbreak, and CKA-Agent. For these methods, individual sub-queries often appear benign and are distributed across multiple turns or agents, rendering defenses that operate on isolated prompts inherently limited. We also observe slight increases in attack success for certain methods, potentially because rephrasing bypasses detection or Llama Guard’s refusals encourage more strategic exploration. Among all methods, Circuit Breaker remains the most effective against CKA-Agent, largely due to its approach of remapping representations associated with harmful processes. Furthermore, Circuit Breaker is notably effective against TAP, aligning with findings in [zou2024improving], likely because TAP relies on surface-level transformations like synonym substitution or role-play rather than obfuscating the underlying harmful intent detected at the representation level. Nevertheless, while representation-level control outperforms pure prompt-level defenses, existing mechanisms remain inadequate for multi-turn scenarios due to their limited ability to aggregate signals and infer intent across conversational contexts. Potential defense directions are further discussed in Sec. 4.5.
@@ 4.3 Alignment Between Human and LLM Judgments
To assess the calibration of our LLM-as-Judge ASR metric, we recruited ten doctoral researchers specializing in Electrical Engineering and Computer Science (EECS) to conduct a human evaluation study. The study employed a between-subjects design on 40 randomly sampled prompt–response pairs (10 per category: FS, PS, V, R), with the evaluation session lasting two hours per annotator. Five annotators evaluated the pairs using only the prompt and response (No Reasoning Provided), while a distinct group of five evaluators assessed the same items with access to the judge model’s reasoning (With Reasoning Provided). As shown in Fig. 7, alignment with the LLM judge increases substantially, from a Spearman correlation of 0.52 in Condition 1 to 0.90 in Condition 2. Crucially, the high variance observed in the No Reasoning setting aligns with our dataset design, which intentionally targets high-stakes domains requiring specialized knowledge (e.g., Chemistry, Biology). Since our evaluators possess deep expertise in EECS but lack specific domain knowledge in these external fields, they exhibited difficulty and a leniency bias when assessing the technical actionability of the responses without guidance. Providing the judge model’s analytical reasoning effectively bridges this domain gap, leading human annotators to evaluations that closely track the LLM’s assessments. These findings confirm that when supported by relevant domain reasoning, the judge model aligns closely with expert human judgment.
@@ 4.5 Evaluating Malicious Intent Detection Capabilities via Single-Session Decomposition
We further posit that CKA-Agent serves as a critical diagnostic framework for studying a fundamental limitation in current alignment: the inability to infer malicious intent when it is semantically distributed across a sequence of innocuous turns. While prior decomposition studies often leverage cross-session attacks to evade history tracking mechanisms, such settings deliberately obscure the model’s inherent multi-turn reasoning capabilities by introducing memory loss as a confounding variable. To rigorously evaluate whether aligned models can detect malicious users when given maximal visibility, we propose the single-session CKA-Agent-Branch as a stress test for intent aggregation. In this setting, each sub-query is submitted alongside the entire conversation history of the current branch. This effectively simulates a context-aware defense where the model has “perfect memory.” Further, we introduce CKA-Agent-Primed, where we explicitly instruct the target model to be aware that a collection of benign sub-queries within the current branch may collectively encode a latent malicious intent. Table 3 validates the utility of this framework. Even with full history visibility in CKA-Agent-Branch, the attack maintains Full Success rates exceeding 92% on Gemini and Claude models. Remarkably, introducing explicit warnings (CKA-Agent-Primed) yields only a modest defense improvement, with success rates dropping by approximately 10% on these models. This limited reduction indicates that heightened awareness alone is insufficient. Furthermore, while GPT-oss exhibits stronger baseline resistance in the Branch setting, the additional explicit prompts provide limited marginal gains, leaving the model vulnerable in over 73% / 82% of cases. This consistently high failure rate across all evaluated models exposes a structural deficiency: current alignment processes fine-tune models to detect atomic harmful prompts, but fail to instill the capability to aggregate intent over extended dialogs. This observation echoes the previous studies [zhou2024speak, liu2024lost, laban2025llms]. Consequently, we argue that CKA-Agent-Branch and CKA-Agent-Primed provide robust testbeds for future research, distinguishing between simple pattern matching and genuine, long-horizon intent understanding.
@@
  CKA-Agent-Primed 0.883 0.086 0.012 0.019 0.852 0.086 0.031 0.031 0.821 0.142 0.000 0.037 0.864 0.099 0.019 0.019
Table 3: Performance comparison between CKA-Agent, CKA-Agent-Branch, and CKA-Agent-Primed across different models and datasets.